OpenStudy (anonymous):

-4y + 2y =-10 as an ordered pair? Can someone help?

OpenStudy (misty1212):

Hi!!

OpenStudy (misty1212):

question doesn't make any sense as it is written

OpenStudy (misty1212):

can you post a screen shot or sommat?

OpenStudy (anonymous):

It's Solving System of Equations by Substitution. I got the substitution part. Because x=-4y so i plugged that in x+2y=-10. So now it would be -4y+2y=-10. I don't know what to do from there.

OpenStudy (misty1212):

ok now it makes sense

OpenStudy (misty1212):

where you see \(x\) in the first equation, replace it by \(-4y\) and solve \[-4y+2y=-10\] for \(y\)

OpenStudy (misty1212):

is the problem solving that equation?

OpenStudy (anonymous):

Yes, i don't know how.